OpenStudy (anonymous):

State domain, range, period, vertical asymptotes, zeros, symmetry and y-intercept of: y = -2tan(3x + 180°) + 3

OpenStudy (anonymous):

@mhmdrz91 help please

OpenStudy (anonymous):

Domain\[{x \in R : 6 (x+60 °)2 \pi n+\pi\ and\ 6 (x+60 °)<\pi (2 n+3)\ and\ n \in Z}\]Range\[R\]and Period is \[\frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

what does the pie symbol stand for?

OpenStudy (anonymous):

vertical asymptote ? @mhmdrz91

OpenStudy (anonymous):

are you there ???

OpenStudy (anonymous):

vertical asymptote is\[(2k−1)\frac{\pi}{6}\ that\ k∈Z\]

OpenStudy (anonymous):

@mhmdrz91 zeros, symmetry and y-intercept of:

OpenStudy (anonymous):

??

OpenStudy (anonymous):

you can just email me what you get